Find the area of the region bounded by the x and y axes and the line y= 2x +5.
Answer by sachi(548)

Find the area of the region bounded by the x and y axes and the line y= 2x +5.
y= 2x +5
now putting x=0 y=5
& putting y=0 x=-5/2
so the line cuts x-axis at (-5/2,0) & y-axis at (0,5)
so area forms by the line a right angled triangle with x& y axes
where the base is 5/2 units & height is 5 units
so the area =1/2*5/2*5=25/4=6.25 sq units